# Field-sensitive dislocation bound states in two-dimensional $d$-wave altermagnets

### v1

@@ III Field-sensitive dislocation Majorana zero modes
This result indicates that, despite the impossibility of chiral Majorana modes when $M_{z}=0$, Majorana zero modes can be created at dislocations with Burgers vectors equal to either ±$\pm\bm{e_{x}}$ or ±$\pm\bm{e_{y}}$. Before explicitly showing this, we first complete the analysis of the bulk topology when $M_{z}\neq 0$.
@@
Figure 4: BdG energy spectrum for a sample of ribbon geometry. The system takes open boundary directions in the $y$ ($x$) direction and periodic boundary conditions in the $x$ ($y$) direction for the top (bottom) row. In (a) and (b), $M_{z}=0$, $C=0$, and the locations of edge-state spectrum crossings suggest $(\nu_{x},\nu_{y})=(1,1)$. In (c) and (d), $M_{z}=-0.3$, $C=-1$, and the locations of edge-state spectrum crossings suggest $(\nu_{x},\nu_{y})=(1,0)$. In (e) and (f), $M_{z}=0.3$, $C=1$, and the locations of edge-state spectrum crossings suggest $(\nu_{x},\nu_{y})=(0,1)$. Common parameters are $t=0.2$, $t_{\rm AM}=0.3$, $\lambda=0.3$, $\Delta_{s}=0.2$, and $\mu=-1.1$.
@@
Similar to the previous nonsuperconducting case, a finite $M_{z}$ breaks the $C_{4z}\mathcal{T}$ symmetry and can induce topological superconducting phases characterized by nonzero Chern numbers. The topological phase diagram can be easily determined since the change of Chern number is associated with the close of bulk energy gapSau et al. (2010), which takes place when $M_{z}=\{\pm\sqrt{(4t\pm\mu)^{2}+\Delta_{s}^{2}}$, ±$\pm 4t_{\rm AM}\pm\sqrt{\mu^{2}+\Delta_{s}^{2}}$}. Here we are not interested in determining the complete phase diagram. Rather, we are interested in the region where the exchange field is weak, i.e., when $M_{z}$ is comparable to $\Delta_{s}$, and when the chemical potential is close to the critical points between the weak topological superconductor and trivial superconductors, i.e., when $\mu$ is close to $\mu_{c,\pm}=\pm\sqrt{16t_{\rm AM}^{2}-\Delta_{s}^{2}}$. Without loss of generality, we take $\mu=\mu_{c,-}+\delta$ where $\delta$ is a small real positive constant. Then, increasing the exchange field’s strength from zero, the energy gap closes at X when $M_{z}=4t_{\rm AM}-\sqrt{\mu^{2}+\Delta_{s}^{2}}\simeq\delta$ (assuming $\Delta_{s}\ll t_{\rm AM}$), and at Y when $M_{z}=-4t_{\rm AM}+\sqrt{\mu^{2}+\Delta_{s}^{2}}\simeq-\delta$. The close of energy gap at X changes the total Chern number of the bands below $E=0$ from $C=0$ to $C=1$, and the weak topological indices $(\nu_{x},\nu_{y})$ from $(1,1)$ to $(0,1)$. In contrast, the close of energy gap at Y changes the total Chern number of the bands below $E=0$ from $C=0$ to $C=-1$, and the weak topological indices $(\nu_{x},\nu_{y})$ from $(1,1)$ to $(1,0)$.
@@
In Fig.4, we show the energy spectrum for a sample of ribbon geometry. In Figs.4(a) and 4(b), we see that there are two counter-propagating gapless states on one edge and the edge-state spectrum crossing occurs at time-reversal invariant momenta on both $x$-normal and $y$-normal edges; this is consistent with $C=0$ and $(\nu_{x},\nu_{y})=(1,1)$ for $M_{z}=0$ and $4t_{\rm AM}>\sqrt{\mu^{2}+\Delta_{s}^{2}}$. In Figs.4(c) and 4(d), the number of chiral edge states and the locations of edge-state spectrum crossings suggest that $C=-1$ and $(\nu_{x},\nu_{y})=(1,0)$, which is also consistent with the values as analyzed above. In Figs.4(e) and 4(f), the results show that a reversal of the exchange field’s direction reverses the chirality of the gapless edge state, and leads to a switching of the locations of edge-state spectrum crossings on the $x$-normal and $y$-normal edges. This again agrees with the sign change of the Chern number and the switching of the values of $\nu_{x}$ and $\nu_{y}$ accompanying this process.
@@
Placing a cross-shaped defect in the superconducting system and diagonalizing the Hamiltonian under periodic boundary conditions in both directions, we find the presence of Majorana zero modes at dislocations when the topological criterion $\textbf{B}\cdot\textbf{G}_{\nu}=\pi\,(\text{mod}\,2\pi)$ is fulfilled. To be specific, when the system is a weak topological superconductor with $(\nu_{x},\nu_{y})=(1,1)$ (indicated by the purple region of Fig.5), we find that there are four Majorana zero modes with their wave functions localized at the cores of the dislocations, as shown in the inset on top of the purple region. The result suggests that dislocations with Burgers vector $\textbf{B}=\pm\bm{e_{x}}$ and ±$\pm\bm{e_{y}}$ all harbor Majorana zero modes at their cores, which agrees with the topological criterion. In comparison, when $\nu_{x}$ ($\nu_{y}$) is trivialized by the exchange field, the two Majorana zero modes at the dislocations with $\textbf{B}=\pm\bm{e_{x}}$ (±$\pm\bm{e_{y}}$) disappear, while the two Majorana zero modes at the dislocations with $\textbf{B}=\pm\bm{e_{y}}$ (±$\pm\bm{e_{x}}$) remain intact; this is clearly shown in the inset on top of the red (blue) region in Fig.5. We can draw an important conclusion from Fig.5. Namely, when the chemical potential is close to $\mu_{c,-}$, tuning the direction of a weak exchange field can control the presence or absence of Majorana zero modes at a dislocation. These results suggest that the unique spin-splitting band structure induced by AM and Rashba SOC provides a basis for the realization of controllable Majorana modes.
@@
Figure 5: Phase diagram around the critical point with $\mu_{c,-}=-\sqrt{16t_{\rm AM}^{2}-\Delta_{s}^{2}}$ and $M_{c}=0$. The phase boundaries are given by $M_{z}=\pm(4t_{\rm AM}-\sqrt{\mu^{2}+\Delta_{s}^{2}})$. The weak topological indices $(\nu_{x},\nu_{y})$ in the blue, white, red and purple regions take values of $(1,0)$, $(0,0)$, $(0,1)$ and (1,1), respectively. The four insets at the corresponding regions show the number and distributions of Majorana zero modes at the two pairs of dislocations with $\textbf{B}=\pm\bm{e_{x}}$ and ±$\pm\bm{e_{y}}$. The corresponding parameters of the insets on the left, right, top and bottom are $(\mu,M_{z})=(-1.183,-0.3)$, $(\mu,M_{z})=(-1.183,0.3)$, $(\mu,M_{z})=(-0.6,0)$ and $(\mu,M_{z})=(-1.4,0)$. Common parameters are $t=0.2$, $t_{\rm AM}=0.3$, $\lambda=0.3$ and $\Delta_{s}=0.2$.
@@ IV Discussions and conclusions
From a symmetry perspective, the change from one set of topological dislocation bound states into their $C_{4z}$-rotational counterparts upon reversing the exchange field’s direction is a direct manifestation of the $C_{4z}\mathcal{T}$ symmetry of the pristine altermagnetic system. In the light of this, an observation of this switching behavior of topological dislocation bound states can be taken as a strong signature of $d$-wave AM. In experiments, the field-sensitive dislocation bound states are robust due to their topological origin and can be easily detected by scanning tunneling microscopy. Given the ubiquitous presense of topological defects in real materials and a continuously accumulated list of candidates for altermagnetic materials, our findings unveil a promising route to an effective experimental diagnosis of AM in these materials.
@@ Acknowledgements
D. Z., D. L., Z.-Y. Z, and Z. Y. are supported by the National Natural Science Foundation of China (Grant No. 12174455), Natural Science Foundation of Guangdong Province (Grant No. 2021B1515020026), and Guangdong Basic and Applied Basic Research Foundation (Grant No. 2023B1515040023). Z. W. is supported by National Key R&\&&D Program of China (Grant No. 2022YFA1404103), NSFC (Grant No. 11974161) and Shenzhen Science and Technology Program (Grant No. KQTD20200820113010023).

### v2

@@ III Field-sensitive dislocation Majorana zero modes
This result indicates that, despite the impossibility of chiral Majorana modes when $M_{z}=0$, Majorana zero modes can be created at dislocations with Burgers vectors equal to either $\pm\bm{e_{x}}$ or $\pm\bm{e_{y}}$. Before explicitly showing this, we first complete the analysis of the bulk topology when $M_{z}\neq 0$.
@@
Figure 4: BdG energy spectrum for a sample of ribbon geometry. The system takes open boundary directions in the $y$ ($x$) direction and periodic boundary conditions in the $x$ ($y$) direction for the top (bottom) row. In (a) and (b), $M_{z}=0$, $C=0$, and the locations of edge-state spectrum crossings suggest $(\nu_{x},\nu_{y})=(1,1)$. In (c) and (d), $M_{z}=-0.3$, C=1C=-1, and the locations of edge-state spectrum crossings suggest $(\nu_{x},\nu_{y})=(1,0)$. In (e) and (f), $M_{z}=0.3$, $C=1$, and the locations of edge-state spectrum crossings suggest $(\nu_{x},\nu_{y})=(0,1)$. Common parameters are $t=0.2$, $t_{\rm AM}=0.3$, $\lambda=0.3$, $\Delta_{s}=0.2$, and $\mu=-1.1$.
@@
Similar to the previous nonsuperconducting case, a finite $M_{z}$ breaks the $C_{4z}\mathcal{T}$ symmetry and can induce topological superconducting phases characterized by nonzero Chern numbers. The topological phase diagram can be easily determined since the change of Chern number is associated with the close of bulk energy gapSau et al. (2010), which takes place when $M_{z}=\{\pm\sqrt{(4t\pm\mu)^{2}+\Delta_{s}^{2}}$, $\pm 4t_{\rm AM}\pm\sqrt{\mu^{2}+\Delta_{s}^{2}}$}. Here we are not interested in determining the complete phase diagram. Rather, we are interested in the region where the exchange field is weak, i.e., when $M_{z}$ is comparable to $\Delta_{s}$, and when the chemical potential is close to the critical points between the weak topological superconductor and trivial superconductors, i.e., when $\mu$ is close to $\mu_{c,\pm}=\pm\sqrt{16t_{\rm AM}^{2}-\Delta_{s}^{2}}$. Without loss of generality, we take $\mu=\mu_{c,-}+\delta$ where $\delta$ is a small real positive constant. Then, increasing the exchange field’s strength from zero, the energy gap closes at X when $M_{z}=4t_{\rm AM}-\sqrt{\mu^{2}+\Delta_{s}^{2}}\simeq\delta$ (assuming $\Delta_{s}\ll t_{\rm AM}$), and at Y when $M_{z}=-4t_{\rm AM}+\sqrt{\mu^{2}+\Delta_{s}^{2}}\simeq-\delta$. The close of energy gap at X changes the total Chern number of the bands below $E=0$ from $C=0$ to $C=1$, and the weak topological indices $(\nu_{x},\nu_{y})$ from $(1,1)$ to $(0,1)$. In contrast, the close of energy gap at Y changes the total Chern number of the bands below $E=0$ from $C=0$ to C=1C=-1, and the weak topological indices $(\nu_{x},\nu_{y})$ from $(1,1)$ to $(1,0)$.
@@
In Fig.4, we show the energy spectrum for a sample of ribbon geometry. In Figs.4(a) and 4(b), we see that there are two counter-propagating gapless states on one edge and the edge-state spectrum crossing occurs at time-reversal invariant momenta on both $x$-normal and $y$-normal edges; this is consistent with $C=0$ and $(\nu_{x},\nu_{y})=(1,1)$ for $M_{z}=0$ and $4t_{\rm AM}>\sqrt{\mu^{2}+\Delta_{s}^{2}}$. In Figs.4(c) and 4(d), the number of chiral edge states and the locations of edge-state spectrum crossings suggest that C=1C=-1 and $(\nu_{x},\nu_{y})=(1,0)$, which is also consistent with the values as analyzed above. In Figs.4(e) and 4(f), the results show that a reversal of the exchange field’s direction reverses the chirality of the gapless edge state, and leads to a switching of the locations of edge-state spectrum crossings on the $x$-normal and $y$-normal edges. This again agrees with the sign change of the Chern number and the switching of the values of $\nu_{x}$ and $\nu_{y}$ accompanying this process.
@@
Placing a cross-shaped defect in the superconducting system and diagonalizing the Hamiltonian under periodic boundary conditions in both directions, we find the presence of Majorana zero modes at dislocations when the topological criterion $\textbf{B}\cdot\textbf{G}_{\nu}=\pi\,(\text{mod}\,2\pi)$ is fulfilled. To be specific, when the system is a weak topological superconductor with $(\nu_{x},\nu_{y})=(1,1)$ (indicated by the purple region of Fig.5), we find that there are four Majorana zero modes with their wave functions localized at the cores of the dislocations, as shown in the inset on top of the purple region. The result suggests that dislocations with Burgers vector $\textbf{B}=\pm\bm{e_{x}}$ and $\pm\bm{e_{y}}$ all harbor Majorana zero modes at their cores, which agrees with the topological criterion. In comparison, when $\nu_{x}$ ($\nu_{y}$) is trivialized by the exchange field, the two Majorana zero modes at the dislocations with $\textbf{B}=\pm\bm{e_{x}}$ ($\pm\bm{e_{y}}$) disappear, while the two Majorana zero modes at the dislocations with $\textbf{B}=\pm\bm{e_{y}}$ ($\pm\bm{e_{x}}$) remain intact; this is clearly shown in the inset on top of the red (blue) region in Fig.5. We can draw an important conclusion from Fig.5. Namely, when the chemical potential is close to $\mu_{c,-}$, tuning the direction of a weak exchange field can control the presence or absence of Majorana zero modes at a dislocation. These results suggest that the unique spin-splitting band structure induced by AM and Rashba SOC provides a basis for the realization of controllable Majorana modes.
@@
Figure 5: Phase diagram around the critical point with $\mu_{c,-}=-\sqrt{16t_{\rm AM}^{2}-\Delta_{s}^{2}}$ and $M_{c}=0$. The phase boundaries are given by $M_{z}=\pm(4t_{\rm AM}-\sqrt{\mu^{2}+\Delta_{s}^{2}})$. The weak topological indices $(\nu_{x},\nu_{y})$ in the blue, white, red and purple regions take values of $(1,0)$, $(0,0)$, $(0,1)$ and (1,1), respectively. The insets corresponding to the four topologically distinct regions show the number and distributions of Majorana zero modes at the two pairs of dislocations with $\textbf{B}=\pm\bm{e_{x}}$ and $\pm\bm{e_{y}}$ in a system with lattice size $N_{x}\times N_{y}=29\times 29$ and periodic boundary conditions in both $x$ and $y$ directions. The parameters of the insets that correspond to the selected point in the left, right, top and bottom regions are $(\mu,M_{z})=(-1.183,-0.3)$, $(\mu,M_{z})=(-1.183,0.3)$, $(\mu,M_{z})=(-0.6,0)$ and $(\mu,M_{z})=(-1.4,0)$, respectively. Common parameters are $t=0.2$, $t_{\rm AM}=0.3$, $\lambda=0.3$ and $\Delta_{s}=0.2$.
@@ IV Discussions and conclusions
From a symmetry perspective, the change from one set of topological dislocation bound states into their $C_{4z}$-rotational counterparts upon reversing the exchange field’s direction is a direct manifestation of the $C_{4z}\mathcal{T}$ symmetry of the pristine altermagnetic system. In the light of this, an observation of this switching behavior of topological dislocation bound states can be taken as a strong signature of $d$-wave AM. In experiments, the field-sensitive dislocation bound states are robust due to their topological origin and can be easily detected by scanning tunneling microscopy. Up to now, a list of 2D materials with $d$-wave AM have been predictedBai et al. , which makes the experimental realization of our proposal quite promising. Particularly, we note that $d$-wave AM has recently predicted to be a possible origin of the time-reversal symmetry breaking observed in FeSe grown on SrTiO3Mazin et al. (2023). Our findings in the current paper may provide a way to test this prediction in this intensively-studied material. To conclude, given the ubiquitous presense of topological defects in real materials and a continuously accumulated list of candidates for altermagnetic materials, our findings unveil a promising route to an effective experimental diagnosis of AM in these materials.
@@ Acknowledgements
D. Z., D. L., Z.-Y. Z, and Z. Y. are supported by the National Natural Science Foundation of China (Grant No. 12174455), Natural Science Foundation of Guangdong Province (Grant No. 2021B1515020026), and Guangdong Basic and Applied Basic Research Foundation (Grant No. 2023B1515040023). Z. W. is supported by National Key R&\&D Program of China (Grant No. 2022YFA1404103), NSFC (Grant No. 11974161) and Shenzhen Science and Technology Program (Grant No. KQTD20200820113010023).